# What Expressivity Theory Misses: Message Passing Complexity for GNNs

### v1

@@ 4.2 MPC describes theoretical and practical limitations of MPNNs
Having defined our complexity measure MPC, we now demonstrate that it unifies theoretical expressivity with practical GNN limitations in a single framework. We first establish that MPC preserves impossibility results from iso expressivity theory, then show how it captures practical phenomena such as over-squashing and under-reaching. We begin by characterizing when MPC becomes infinite, which precisely corresponds to tasks that are theoretically impossible for a given architecture $\mathcal{M}$}. [^1] We defer all proofs to App. C. {restatable}[Infeasibility]theoremInfeasibility The complexity for ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$ and function $f$ is ${\textup{MPC}}_{\mathcal{M}}(f_{v},{G})=\infty$ if and only if there exist ${G}^{\prime},w\in{{\mathcal{G}}_{\mathcal{V}}}$ such that $f_{v}({G})\neq f_{w}({G}^{\prime})$ but ${M}_{v}({G})={M}_{w}({G}^{\prime})$ for all model instantiations ${M}\in{\mathcal{M}}$. This shows that MPC becomes infinite precisely when a model architecture fundamentally cannot distinguish between two different cases that require different outputs. We can also globally characterize the functions an architecture can express: {restatable}lemmarecover There exists no model instantiation ${M}\in{\mathcal{M}}$ such that ${M}_{v}({G})=f_{v}({G})$ for all ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$ if and only if there exists ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$ with ${\textup{MPC}}_{\mathcal{M}}(f_{v},{G})=\infty$. By choosing an isomorphism test (like the WL test) as task $f$, we can recover all impossibility statements from iso expressivity theory (Lemma˜C.7). Thus, MPC subsumes iso expressivity theory while providing a continuous difficulty measure for possible tasks. Having established when MPC becomes infinite, we now characterize how MPC scales with function granularity: a finer, more discriminative task can never have lower complexity than the coarser task it refines. For example, counting the exact number of cycles in a graph should be more complex than merely detecting whether any cycle exists. {restatable}[Function refinement]theoremrefinement Let $f$ be a function that is more fine-grained than $g$, i.e., $f\vDash_{{\mathcal{G}}_{\mathcal{V}}}g$. Then for any ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$: ${\textup{MPC}}_{\mathcal{M}}(f_{v},G)\geq{\textup{MPC}}_{\mathcal{M}}(g_{v},G)$. Like other complexity measures, MPC satisfies compositionality: solving tasks jointly cannot be more complex than solving them separately, and may be easier when tasks share information. {restatable}[Task Triangle Inequality]lemmatriangle Let $f$ and $g$ be functions, and $\|$ denote concatenating function outputs. Then for any ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$: ${\textup{MPC}}_{\mathcal{M}}(f_{v}\|g_{v},{G})\leq{\textup{MPC}}_{\mathcal{M}}(f_{v},{G})+{\textup{MPC}}_{\mathcal{M}}(g_{v},{G}).$ However, MPC connects not only to properties of classical complexity and expressivity theory but also captures real-world limitations of MPNNs such as over-squashing and under-reaching. These phenomena, extensively studied as significant constraints on MPNN performance, are overlooked by traditional expressivity theory. We first relate MPC to over-squashing by showing that it is lower bounded in terms of the $L$-step random walk probability from $v$ to $u$, $({{\bm{I}}}^{L})_{vu}$, a quantity that also motivated our design of lossyWL in §˜4.1.
@@
A lower probability $({{\bm{I}}}^{L})_{vu}$ indicates that node $v$ receives less gradient signal from $u$, making it more susceptible to over-squashing Giovanni et al. [2023], Xu et al. [2018]. Our complexity measure captures this: when $({{\bm{I}}}^{L})_{vu}$ is small, tasks requiring information flow from $u$ to $v$ have high complexity. As a simple special case, we consider under-reaching, which occurs when the number of MPNN layers is less than the graph diameter, preventing nodes from receiving information from distant parts of the graph Barceló et al. [2020], Alon and Yahav [2021]:
@@
These results show that MPC captures both theoretical expressivity and practical limitations of GNNs, providing a more unified framework that bridges these previously separate perspectives.
@@ 5 MPC in Practice: Explaining Empirical MPNN Behavior
We now demonstrate that MPC quantitatively explains empirical MPNN behavior. Through analysis of fundamental graph tasks—retaining information, propagating information, and extracting topological features—we show that: 1) MPC enables the derivation of meaningful task-specific complexity bounds; 2) trends in MPC complexity correlate strongly with empirical performance, better reflecting the practical behavior of MPNNs than iso expressivity theory; 3) optimal architectures vary by task, with success determined by minimizing task-specific MPC complexity through appropriate inductive biases, not maximizing expressivity. Experimental Setup: We verify this for a wide range of architectures, including an MLP baseline (message passing on an empty graph), standard models like GCN Kipf and Welling [2017], GIN Xu et al. [2019], GraphSage Hamilton et al. [2017], and a GCN with a virtual node, GCN-VN Gilmer et al. [2017], as well as higher-order models like GSN Bouritsas et al. [2023] which incorporates substructure-information as node-feature, FragNet Wollschläger et al. [2024] which builds a higher-level graph of fragments and the topological-inspired model CIN Bodnar et al. [2021a]. We evaluate on random $r$-regular graphs $\mathcal{G}$} and show that the results transfer to graphs from the ZINC dataset Dwivedi et al. [2023] and the long-range graph benchmark Dwivedi et al. [2022]. We provide theoretical complexity bounds and Monte-Carlo simulated complexity values. We compare MPC to the WL-based WLC baseline rather than directly to other expressivity measures, as they have fundamental limitations for task-specific difficulty analysis: they typically provide only global, task-agnostic architecture rankings Zhang et al. [2024, 2023] or impose restrictive assumptions incompatible with our tasks Giovanni et al. [2024] (see §˜7).
Retaining information First, we evaluate the task $f_{v}({G})={\bm{X}}_{v}$, which tests a model’s ability to retain its initial node features. This task is fundamentally important across all domains with informative node features (e.g., atomic type in molecules). Although seemingly simple, it is closely linked to the well-studied over-smoothing Li et al. [2018] phenomenon, where deeper GNNs lose distinguishability between nodes as their representations converge, making it difficult to recover the original node features Giraldo et al. [2023]. Our complexity measure captures this difficulty as MPC increases at least linearly with depth $L$: {restatable}lemmaretaining Assume the degree $r\geq 2$. Then the expected MPC complexity for this task $\mathbb{E}_{{G},v\sim{{\mathcal{G}}_{\mathcal{V}}}}[{\textup{MPC}}_{\mathcal{M}}(f_{v},G)]$ grows at least linearly with $L$, i.e., is in $\Omega(L)$, for all MPNNs $\mathcal{M}$} .
@@
For empirical validation, we train all model architectures on 2000 randomly generated $3$-regular graphs with varying numbers of layers $L$. As shown in Figs.˜3 and 9, the complexity measure based on iso expressivity theory, WLC, assigns constant zero complexity regardless of depth $L$, indicating only theoretical solvability. In contrast, MPC shows perfect negative Spearman correlation with accuracy for most architectures, capturing the increasing difficulty with $L$. Only CIN and GraphSage maintain perfect accuracy throughout, due to explicit residual connection optimization—an implementation detail our framework abstracts from. These results demonstrate that, unlike binary expressivity measures, MPC accurately quantifies the progressive difficulty of this task for most architectures, capturing important real-world limitations beyond theoretical impossibility statements.
@@
In summary, these results demonstrate why MPC offers a more complete and practical understanding of MPNN capabilities than classical expressivity theory. Unlike binary expressivity tests, MPC explains why even highly expressive architectures, such as GSN, FragNet, and CIN, struggle with fundamental tasks like retaining or propagating node information—empirically validating our critique in §˜3 that iso expressivity poorly captures practical capabilities. Furthermore, MPC accounts for the success of architectural choices like virtual nodes, which consistently improve performance on real-world long-range tasks Cai et al. [2023] despite unchanged iso expressivity Wollschläger et al. [2024].
@@
$\text{For CIN \& FragNet:}\>\>O(\log(sr))\quad\quad\text{For GSN:}\>\>\leq s/2\log(r+1)\quad\quad\text{For all others:}\>\>\geq s\log(r)$
Figure 6: Test average precision (AP) for the ring transfer task compared with complexity values MPC and WLC (MPC values for SAGE & GCN are lower bounds). MPC (in contrast to WLC) is negatively Spearman correlated with AP and captures the advantage of cycle-oriented inductive biases in GSN and FragNet. All results in Fig.˜18, for ZINC in Fig.˜23.
These bounds reveal that MPC captures crucial architectural differences missed by iso expressivity: standard MPNNs have complexity scaling linearly with cycle size, while CIN and FragNet achieve logarithmic scaling, capturing their cycle-oriented inductive biases. Experiments on random 4-regular graphs confirm these theoretical predictions, with trends in MPC complexity matching empirical performance (see Fig.˜6, in relation to dataset size: Fig.˜19).
Because of the unique node labels, all (at least WL expressive) architectures are theoretically able to solve this task, yet performance varies dramatically. While iso expressivity theory cannot capture these differences, MPC provides a clear explanation: GSN, CIN and FragNet’s architectural bias towards cycles reduces their message-passing complexity for this task, providing them with a performance advantage. This insight extends to real-world graphs from the ZINC molecular regression benchmark Dwivedi et al. [2023] (Fig.˜6), where the identification of molecular rings is crucial. As shown in Tab.˜2, architectures with lower ring detection complexity consistently achieve better performance on ZINC on both synthetic and real tasks. The superior performance of higher-order MPNNs like GSN, FragNet or CIN cannot be explained by their greater iso expressivity—as discussed in §˜3, almost all graphs in ZINC are already WL-distinguishable—but by their cycle-oriented inductive biases that reduce the complexity of learning molecular structures.
@@ 6 Limitations & Future Work
While we have demonstrated that our complexity measure closely aligns with empirical performance for many tasks, it abstracts away from the parametrization and implementation of the upd and agg functions and the optimization process as discussed in §˜5. Additionally, our MPNN framework does not include models with attention. However, we repeat all our experiments with the most common attention-based MPNNs in §˜D.1, showing similar empirical trends to the MPNNs that we consider. Secondly, exact target functions in real-world applications are usually unknown. Still, we can analyze two types of proxy tasks: fundamental capabilities required across all domains (like retaining information) and domain-specific operations (like cycle detection). As shown in Tab.˜2, the insights from these proxy tasks can translate to real-world performance. Third, exact complexity values can be difficult to compute for sophisticated tasks. However, as our experiments demonstrate, theoretical bounds often provide sufficient practical insight. Finally, the complexity values should not be interpreted in isolation to determine whether an architecture can solve a given task, as they only characterize the message-passing complexity of a task. Instead, they are most useful for identifying trends and comparing architectures across a single dimension of freedom.
@@
More broadly, we hope that MPC will enable more principled architectural design by providing fine-grained analysis of complexity for specific capabilities, offering insights more closely aligned with empirical performance than traditional expressivity theory. Rather than pursuing maximally expressive models, MPC can reveal practical limitations of current architectures and guide future research toward architectures with low complexity for specific capabilities relevant to target domains. As a concrete example, consider substructure identification. Our analysis in §˜5 indicates that all considered MPNNs without additional pre-processing steps exhibit very high complexity for this task, even when theoretically solvable. This suggests that architectural modifications alone may be insufficient, pointing toward studying expressive preprocessing steps or positional encodings —as successfully employed by FragNet Wollschläger et al. [2024] and CIN Bodnar et al. [2021a]— that directly reduce task complexity. More generally, this exemplifies how MPC can shift architectural design from maximizing universal expressivity toward strategically minimizing complexity for domain-specific requirements.
@@ 7 Related Work
Iso-expressivity frameworks. The predominant approach for analyzing GNNs compares their ability to distinguish non-isomorphic graphs relative to the WL test and its extensions Morris et al. [2019]. This has motivated extensive research on developing higher-order architectures that surpass standard MPNNs in expressivity Morris et al. [2021], Bodnar et al. [2021a], Giusti et al. [2023], Wollschläger et al. [2024]. However, as discussed in §˜3, iso-expressivity theory relies on idealized assumptions and provides only binary characterizations, limiting its practical relevance.
Alternative expressivity characterizations. Beyond iso-expressivity, several frameworks characterize MPNN capabilities through specific tasks. Zhang et al. [2024] rank architectures by the set of substructures they can recognize, while Zhang et al. [2023] compare models through their ability to solve graph biconnectivity. Wang and Zhang [2022] analyze spectral MPNNs through their ability to learn polynomial filters. Logic-based approaches Barceló et al. [2020], Grohe [2024] characterize learnable functions through fragments of first-order logic, providing more nuanced insights by considering effects of different activation functions Khalife and Tonelli-Cueto [2024]. However, these expressivity characterizations share key limitations with iso expressivity: they remain binary (can/cannot solve) and assume lossless information propagation, limiting their insights for real-world MPNNs.
Continuous graph distances. To move beyond binary expressivity, some works have proposed continuous graph similarity measures, including tree-based Chuang and Jegelka [2022], graphon-based Böker et al. [2023], and Wasserstein-based distances Chen et al. [2022]. However, these metrics are task-agnostic and architecture-independent, limiting their relevance for explaining practical GNN performance.
Practical GNN limitations. A parallel line of research studies individual practical limitations of GNNs, such as over-smoothing Li et al. [2018], under-reaching Barceló et al. [2020], and over-squashing Alon and Yahav [2021], Giovanni et al. [2023], Topping et al. [2022]. Most relevant to our approach is the work by Giovanni et al. [2024], which derives expressivity limitations from over-squashing theory. They show tasks become impossible when the required "mixing" between nodes (measured via the Hessian) exceeds what MPNNs can generate. Like MPC, this can capture practical impossibilities arising from over-squashing. However, their framework considers only pairwise interactions with restrictive assumptions on the task (twice differentiable tasks, not dependent on graph topology).

### v2

@@ 4.2 MPC describes theoretical and practical limitations of MPNNs
Having defined our complexity measure MPC, we now demonstrate that it unifies theoretical expressivity with practical GNN limitations in a single framework. We first establish that MPC preserves impossibility results from iso expressivity theory, then show how it captures practical phenomena such as over-squashing and under-reaching. We begin by characterizing when MPC becomes infinite, which precisely corresponds to tasks that are theoretically impossible for a given architecture $\mathcal{M}$}. [^2] We defer all proofs to App. C. {restatable}[Infeasibility]theoremInfeasibility The complexity for ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$ and function $f$ is ${\textup{MPC}}_{\mathcal{M}}(f_{v},{G})=\infty$ if and only if there exist ${G}^{\prime},w\in{{\mathcal{G}}_{\mathcal{V}}}$ such that $f_{v}({G})\neq f_{w}({G}^{\prime})$ but ${M}_{v}({G})={M}_{w}({G}^{\prime})$ for all model instantiations ${M}\in{\mathcal{M}}$. This shows that MPC becomes infinite precisely when a model architecture fundamentally cannot distinguish between two different nodes that require different outputs. We can also globally characterize the functions an architecture can express: {restatable}lemmarecover There exists no model instantiation ${M}\in{\mathcal{M}}$ such that ${M}_{v}({G})=f_{v}({G})$ for all ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$ if and only if there exists ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$ with ${\textup{MPC}}_{\mathcal{M}}(f_{v},{G})=\infty$. By choosing an isomorphism test (like the WL test) as task $f$, we can recover all impossibility statements from iso expressivity theory (Lemma˜C.7). Thus, MPC subsumes iso expressivity theory while providing a continuous difficulty measure for possible tasks. Having established when MPC becomes infinite, we now characterize how MPC scales with function granularity: a finer, more discriminative task can never have lower complexity than the coarser task it refines. For example, counting the exact number of cycles in a graph should be more complex than merely detecting whether any cycle exists. {restatable}[Function refinement]theoremrefinement Let $f$ be a function that is more fine-grained than $g$, i.e., $f\vDash_{{\mathcal{G}}_{\mathcal{V}}}g$. Then for any ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$: ${\textup{MPC}}_{\mathcal{M}}(f_{v},G)\geq{\textup{MPC}}_{\mathcal{M}}(g_{v},G)$. Like other complexity measures, MPC satisfies compositionality: solving tasks jointly cannot be more complex than solving them separately, and may be easier when tasks share information. {restatable}[Task Triangle Inequality]lemmatriangle Let $f$ and $g$ be functions, and $\|$ denote concatenating function outputs. Then for any ${G},v\in{{\mathcal{G}}_{\mathcal{V}}}$: ${\textup{MPC}}_{\mathcal{M}}(f_{v}\|g_{v},{G})\leq{\textup{MPC}}_{\mathcal{M}}(f_{v},{G})+{\textup{MPC}}_{\mathcal{M}}(g_{v},{G}).$ However, MPC connects not only to properties of classical complexity and expressivity theory but also captures real-world limitations of MPNNs such as over-squashing and under-reaching. These phenomena, extensively studied as significant constraints on MPNN performance, are overlooked by traditional expressivity theory. We first relate MPC to over-squashing by showing that it is lower bounded in terms of the $L$-step random walk probability from $v$ to $u$, $({{\bm{I}}}^{L})_{vu}$, a quantity that also motivated our design of lossyWL in §˜4.1.
@@
A lower probability $({{\bm{I}}}^{L})_{vu}$ indicates that node $v$ receives less gradient signal from $u$, making it more susceptible to over-squashing [19, 47]. Our complexity measure captures this: when $({{\bm{I}}}^{L})_{vu}$ is small, tasks requiring information flow from $u$ to $v$ have high complexity. As a simple special case, we consider under-reaching, which occurs when the number of MPNN layers is less than the graph diameter, preventing nodes from receiving information from distant parts of the graph [3, 1]:
@@
These results show that MPC captures both theoretical expressivity and practical limitations of GNNs, providing a more unified framework that bridges both perspectives.
@@ 5 MPC in Practice: Explaining Empirical MPNN Behavior
We now demonstrate that MPC quantitatively explains empirical MPNN behavior. Through analysis of fundamental graph tasks—retaining information, propagating information, and extracting topological features—we show that: 1) MPC enables the derivation of meaningful task-specific complexity bounds; 2) trends in MPC complexity correlate strongly with empirical performance, better reflecting the practical behavior of MPNNs than iso expressivity theory; 3) optimal architectures vary by task, with success determined by minimizing task-specific MPC complexity through appropriate inductive biases, not maximizing expressivity. Experimental Setup: We verify this for a wide range of architectures, including an MLP baseline (message passing on an empty graph), standard models like GCN [31], GIN [48], GraphSage [25], and a GCN with a virtual node, GCN-VN [18], as well as higher-order models like GSN [8] which incorporates substructure-information as node-feature, FragNet [46] which builds a higher-level graph of fragments and the topological-inspired model CIN [5]. We evaluate on random $r$-regular graphs $\mathcal{G}$} and show that the results transfer to graphs from the ZINC dataset [15] and the long-range graph benchmark [14]. We provide theoretical complexity bounds and Monte-Carlo simulated complexity values. We compare MPC to the WL-based WLC baseline rather than directly to other expressivity measures, as they have fundamental limitations for task-specific difficulty analysis: they typically provide only global, task-agnostic architecture rankings [53, 52] or impose restrictive assumptions incompatible with our tasks [20] (see §˜7).
Retaining information First, we evaluate the task $f_{v}({G})={\bm{X}}_{v}$, which tests a model’s ability to retain its initial node features. This task is fundamentally important across all domains with informative node features (e.g., atom type in molecules). Although seemingly simple, it is closely linked to the well-studied over-smoothing [33] phenomenon, where deeper GNNs lose distinguishability between nodes as their representations converge, making it difficult to recover the original node features [21]. Our complexity measure captures this difficulty as MPC increases at least linearly with depth $L$: {restatable}lemmaretaining Assume the degree $r\geq 2$. Then the expected MPC complexity for this task $\mathbb{E}_{{G},v\sim{{\mathcal{G}}_{\mathcal{V}}}}[{\textup{MPC}}_{\mathcal{M}}(f_{v},G)]$ grows at least linearly with $L$, i.e., is in $\Omega(L)$, for all MPNNs $\mathcal{M}$} .
@@
For empirical validation, we train all model architectures on 2000 randomly generated $3$-regular graphs with varying numbers of layers $L$. As shown in Figs.˜3 and 9, the complexity measure based on iso expressivity theory, WLC, assigns constant zero complexity regardless of depth $L$, indicating only theoretical solvability. In contrast, MPC shows perfect negative Spearman correlation with accuracy for most architectures, capturing the increasing difficulty with $L$. Only CIN and GraphSage maintain perfect accuracy throughout, due to explicit residual connection optimization—an implementation choice our framework abstracts from. These results demonstrate that, unlike binary expressivity measures, MPC accurately quantifies the progressive difficulty of this task for most architectures, capturing important real-world limitations beyond theoretical impossibility statements.
@@
In summary, these results demonstrate why MPC offers a more complete and practical understanding of MPNN capabilities than classical expressivity theory. Unlike binary expressivity tests, MPC explains why even highly expressive architectures, such as GSN, FragNet, and CIN, struggle with fundamental tasks like retaining or propagating node information—empirically validating our critique in §˜3 that iso expressivity poorly captures practical capabilities. Furthermore, MPC accounts for the success of architectural choices like virtual nodes, which consistently improve performance on real-world long-range tasks [9] despite unchanged iso expressivity [46].
@@
$\text{For CIN \& FragNet:}\>\>O(\log(sr))\quad\quad\text{For GSN:}\>\>\leq\lceil s/2\rceil\log(r+1)\quad\quad\text{For all others:}\>\>\geq s\log(r)$
Figure 6: Average precision (AP) for the ring task compared with complexity values MPC and WLC (MPC values for SAGE & GCN are lower bounds). MPC (in contrast to WLC) is negatively Spearman correlated with AP and captures the advantage of cycle-oriented inductive biases in GSN and FragNet. See also Figs.˜18 and 23.
These bounds reveal that MPC captures crucial architectural differences missed by iso expressivity: standard MPNNs have complexity scaling linearly with cycle size, while CIN and FragNet achieve logarithmic scaling, capturing their cycle-oriented inductive biases. Experiments on random 4-regular graphs confirm these theoretical predictions, with trends in MPC complexity matching empirical performance (see Figs.˜6 and 18; in relation to dataset size: Fig.˜19).
Because of the unique node labels, all (at least WL expressive) architectures are theoretically able to solve this task, yet performance varies dramatically. While iso expressivity theory cannot capture these differences, MPC provides a clear explanation: GSN, CIN and FragNet’s architectural bias towards cycles reduces their message-passing complexity for this task, providing them with a performance advantage. This insight extends to real-world graphs from the ZINC molecular regression benchmark [15] (Fig.˜6), where the identification of molecular rings is crucial. As shown in Tab.˜2, architectures with lower ring detection complexity consistently achieve better performance on ZINC on both the synthetic and the real task. The superior performance of higher-order MPNNs like GSN, FragNet or CIN cannot be explained by their greater iso expressivity—as discussed in §˜3, almost all graphs in ZINC are already WL-distinguishable—but by their cycle-oriented inductive biases that reduce the complexity of learning molecular structures.
@@ 6 Limitations & Future Work
While we have demonstrated that our complexity measure closely aligns with empirical performance for many tasks, it abstracts away from the parametrization and implementation of the upd and agg functions and the optimization process as discussed in §˜5. Additionally, our MPNN framework does not include models with attention. However, we repeat all our experiments with the most common attention-based MPNNs in §˜D.1, showing similar empirical trends to the MPNNs that we consider. Secondly, exact target functions in real-world applications are usually unknown. Still, we can analyze two types of proxy tasks: fundamental capabilities required across all domains (like retaining information) and domain-specific operations (like cycle detection). As shown in Tab.˜2, the insights from these proxy tasks can translate to real-world performance. Third, exact complexity values can be difficult to compute for sophisticated tasks. However, as our experiments demonstrate, theoretical bounds can provide valuable practical insights. Finally, the complexity values should not be interpreted in isolation to determine whether an architecture can solve a given task, as they only characterize the message-passing complexity of a task. Instead, they are most useful for identifying trends and comparing architectures across a single dimension of freedom.
@@
More broadly, we hope that MPC will enable more principled architectural design: By quantifying fine-grained task complexities, MPC can reveal practical limitations of current architectures and guide the development of models with low complexity for specific capabilities relevant to target domains. As a concrete example, consider substructure identification. Our analysis in §˜5 shows that all considered MPNNs without additional pre-processing steps exhibit very high complexity for this task, even when it is theoretically solvable. This suggests that architectural modifications alone may be insufficient, pointing toward studying expressive preprocessing steps or positional encodings —as successfully employed by FragNet [46] and CIN [5]— that directly reduce task complexity. More generally, this exemplifies how MPC can shift architectural design from maximizing universal expressivity toward strategically minimizing complexity for domain-specific requirements.
@@ 7 Related Work
Iso-expressivity frameworks. The predominant approach for analyzing GNNs compares their ability to distinguish non-isomorphic graphs relative to the WL test and its extensions [36, 48, 37]. This has motivated extensive research on developing higher-order architectures that surpass standard MPNNs in expressivity [37, 5, 22, 46]. However, as discussed in §˜3, iso-expressivity theory relies on idealized assumptions and provides only binary characterizations, limiting its practical relevance.
Alternative expressivity characterizations. Beyond iso-expressivity, several frameworks theoretically characterize MPNN capabilities through specific tasks. Zhang et al. [53] rank architectures by the set of substructures they can recognize, while Zhang et al. [52] compare models through their ability to solve graph biconnectivity. Wang and Zhang [44] analyze spectral MPNNs through their ability to learn polynomial filters. Logic-based approaches [3, 23] characterize learnable functions through fragments of first-order logic, providing more nuanced insights by considering, for example, the effects of different activation functions [29]. However, these expressivity characterizations share key limitations with iso expressivity: they remain binary (can/cannot solve) and assume lossless information propagation, limiting their insights for real-world MPNNs.
Continuous graph distances. To move beyond binary expressivity, some works have proposed continuous graph similarity measures, including tree-based [12], graphon-based [7], and Wasserstein-based distances [10]. However, these metrics are task-agnostic and architecture-independent, limiting their relevance for explaining practical GNN performance.
Practical GNN limitations. A parallel line of research studies individual practical limitations of GNNs, such as over-smoothing [33], under-reaching [3], and over-squashing [1, 19, 41]. Most relevant to our approach is the work by Giovanni et al. [20], which derives expressivity limitations from over-squashing theory. They show tasks become impossible when the required "mixing" between nodes (measured via the Hessian) exceeds what MPNNs can generate. Like MPC, this can capture practical impossibilities arising from over-squashing. However, their framework considers only pairwise interactions with restrictive assumptions on the task (twice differentiable tasks, not dependent on graph topology).